The volume of a cone shaped hole is 48 on ft cubes if the hole is 4 feet deep is hat is the radius of the hole?


1
Expert's answer
2021-03-18T03:27:11-0400

Volume =B.A*height =1/3"\Pi"r2 h

Thus Volume =48ft3 and Height= 4ft

Replacing in the equation

48=1/3"\Pi"r2 4 divide both sides by 4

12=1/3"\Pi"r2 multiply both sides by 3 to remove fraction

36="\Pi"r2 Divide through by "\Pi"

11.46=r2 into 2d.p.

find the square root both sides

3.39=r into 2d.p.

Therefore, radius r=3.39